# Second-Order Constrained Dynamic Optimization

### v2

@@ Abstract
This paper provides an overview, analysis, and comparison of second-order dynamic optimization algorithms, i.e., constrained Differential Dynamic Programming (DDP) and Sequential Quadratic Programming (SQP). Although a variety of these algorithms has been proposed and used successfully, there exists a gap in understanding the key differences and advantages, which we aim to provide in this work. For constrained DDP, we choose methods that incorporate nonlinear programming techniques to handle state and control constraints, including Augmented Lagrangian (AL), Interior Point, Primal Dual Augmented Lagrangian (PDAL), and Alternating Direction Method of Multipliers. Both DDP and SQP are provided in single- and multiple-shooting formulations, where constraints that arise from dynamics are encoded implicitly and explicitly, respectively. As a byproduct of the review, we also propose a single-shooting PDAL DDP which is robust to the growth of penalty parameters and performs better than the normal AL variant. We perform extensive numerical experiments on a variety of systems with increasing complexity to investigate the quality of the solutions, the levels of constraint violation, iterations for convergence, and the sensitivity of final solutions with respect to initialization. The results show that DDP often has the advantage of finding better local minima, while SQP tends to achieve better constraint satisfaction. For multiple-shooting formulation, both DDP and SQP can enjoy informed initial guesses, while the latter appears to be more advantageous in complex systems. It is also worth highlighting that DDP provides favorable computational complexity and feedback gains as a byproduct of optimization.
@@ 1.2 Differential Dynamic Programming
Another approach to handling constraints is applying Nonlinear Programming (NLP) techniques to DDP. This includes the penalty barrier (Frisch, 1955; Fiacco and McCormick., 1968), Augmented Lagrangian (AL) (D., 1969; Hestenes, 1969), Interior Point (IP) (Byrd et al., 1999; Wächter and Biegler, 2005, 2006), Primal Dual Augmented Lagrangian (PDAL) (Gill and Robinson, 2012; Robinson, 2007), and Alternating Direction Method of Multipliers (ADMM) (Boyd et al., 2011).
The combination of DDP and the $\log$ barrier penalty method was proposed in Grandia et al. (2019), where the barrier function is relaxed to allow constraint violations. The penalty function’s coefficient, known as a penalty parameter, is driven to zero in the original formulation but is fixed in this DDP approach. Almubarak et al. (2022) uses a similar approach with an exact barrier function. Despite these approximations, they are shown to work sufficiently well in practice. AL DDP is the most widely used among these combinations of NLP with DDP for inequality constraints (Plancher et al., 2017; Pellegrini and Russell, 2020) and equality constraints (Kazdadi et al., 2021). The method is quite robust in terms of cost reduction, but it may violate constraints, as they are encoded as a part of the cost. Moreover, it may take many iterations to achieve strict feasibility. To alleviate this problem, the researchers proposed extensions in which algorithms switch from AL DDP to other methods when a trajectory approaches convergence (Lantoine and Russell, 2008; Howell et al., 2019; Aoyama et al., 2021). IP DDP (Pavlov et al., 2021) optimizes control variables, Lagrangian multipliers, and slack variables as the original IP method using DDP. PDAL DDP (Jallet et al., 2022a) was proposed most recently. This method is similar to the AL DDP, but the Lagrangian multipliers are also optimized using DDP in contrast to the AL variant. Both AL and PDAL use penalty parameters that are increased during optimization. In AL, large penalty parameters are known to interrupt optimization, whereas PDAL is more robust to them as described in Robinson (2007). The PDAL DDP is presented in multiple-shooting formulation, which we elaborate on in the next paragraph. Using the Alternating Direction Method of Multipliers (ADMM), several variations of constrained DDP have been presented such as in Sindhwani et al. (2017); Zhou and Zhao (2020), which split the problem into smaller subproblems that are solved sequentially. Distributed ADMM-based DDP algorithms have also been proposed for handling constrained multi-agent control problems by utilizing the parallelizable nature of ADMM (Saravanos et al., 2023; Huang et al., 2023).
@@ 1.3 Sequential Quadratic Programming
SQP was introduced as a constrained optimization technique in Wilson (1963), showing its power in many fields including robotics (Kuindersma et al., 2016; Posa et al., 2014), aerospace (Kenway and Martins, 2014; Kamyar and Taheri, 2014) and chemical engineering (Lucia et al., 2013). The approach relies on sequentially solving QP subproblems with quadratically approximated objectives and linearized constraints, generating new nominal trajectories. For dynamical systems, SQP encodes the constraints that arise from linearized dynamics at every time step as equality constraints. SQP also has single- (Diehl et al., 2002), which is also known as condensing, and multiple-shooting formulations. The single-shooting variant eliminates the state variables and only treats the control as decision variables, implicitly satisfying the dynamics. There exists a significant amount of SQP variants that have been proposed in the literature recently for nonlinear optimal control and model predictive control (MPC) (Kouzoupis et al., 2018). One of the main difficulties of the SQP approach was its computational demands to solve QP subproblems. To alleviate this problem Pantoja and Mayne (1991) proposed a stage-wise version of SQP under control inequality constraints. In the same spirit of DDP, this approach solves smaller subproblems at every time step, rather than solving a problem for a whole time horizon. Another approach for improving the time complexity is to exploit the special sparse structure of the constraints arising from dynamics while solving QP. The growth of computational complexity with respect to time horizon can be reduced from cubic to linear, which is as good as DDP. This approach was proposed with active-set (Glad and Jonson, 1984), IP (Steinbach, 1994), and barrier (Wang and Boyd, 2010) methods. The same reduction is achieved by Riccati-based recursion for Linear Quadratic Regulator (LQR) problems (Dohrmann and Robinett, 1997; Rao et al., 1998; Jørgensen et al., 2004). Note that this technique is available only for multiple-shooting SQP because the sparse structure is lost by condensing (Diehl et al., 2009). In the single-shooting variant, rather than working on the complexity, Singh et al. (2022) incorporates the DDP-style closed loop rollout, achieving a faster convergence speed. Recently, Borras et al. (2023) proposed an approach similar to the multiple-shooting SQP, showing its superiority over single- and multiple-shooting DDPs (Mastalli et al., 2020) in tasks without state constraints. This work uses (LQR) and ADMM to solve QP subproblems efficiently and to enforce constraints. Inspired by Stellato et al. (2020), the matrices used in QP are not updated every iteration to improve computational efficiency.
@@ 1.4 Contribution
Although DDP and SQP have a wealth of literature separately, there exist only few works comparing the two approaches. In the unconstrained case - where SQP coincides with Newton’s method with line search - the early works by Murray and Shoemaker (1992); Murray and Yakowitz (1984) had made a comparison of the two approaches. DDP was noted to have the same quadratic convergence properties as Newton’s method, but with the advantage of solving linear equations whose size remains constant w.r.t. the time horizon. However, given the variety of constrained DDP methods available, a thorough comparison of how these methods compare with SQP-based methods is still missing. A comparison between AL DDP and an SQP-based solver SNOPT (Gill et al., 2002) has been made in Howell et al. (2019), where SQP was shown to converge slower in wall clock time. Sindhwani et al. (2017), compared ADMM DDP with control-limited DDP Tassa et al. (2014) and SQP under only control constraints. Xie et al. (2017) also compared their constrained DDP and SQP (SNOPT) using state and control constrained robotic tasks under a time budget, showing the superiority of DDP. However, the details of SQP, including single- or multiple-shooting etc., are not presented.
Towards the aim of bridging the gap between the availability of various algorithms and their understanding, we compare these algorithms from derivations to performance in this paper. In addition, we propose single-shooting PDAL DDP that inherits the merit of PDAL over AL, that is, the robustness to the large penalty parameter, and add it to the comparison.
Our thorough numerical experiments with simple to complex dynamical systems reveal the difference of the algorithms in terms of the quality of solutions, the levels of constraint violation, iterations for convergence, and the sensitivity of the final solutions with respect to initialization. The results indicate that DDP frequently shows its capability to find better local minima, whereas SQP generally performs better in satisfying constraints. It is also shown that although both DDP and SQP have multiple-shooting formulations that can enjoy informed initial guesses, the SQP variant tends to work more reliably in complex systems. Furthermore, our analysis shows that single-shooting DDPs can offer efficient computational complexity, especially when the system is underactuated and the problem has a long time horizon. This computational efficiency is one of the main motivations for users to choose DDP in addition to the feedback gains that DDP provides.
@@
Feedback Gain -
Converges to Optimal -
@@
    Propsed in this paper.
@@
    Originally $N^{3}(n+m)^{3}$, but can be reduced to $N(n+m)^{3}$ with the sparse structure as mentioned in section 5.2.
@@
    Single-shooting SQP cannot enjoy the reduction available in the multiple-shooting variant and has cubic in $N$.
@@ 2.2 Log Barrier method
2.3 Augmented Lagrangian method
@@
2.4 Interior Point method
@@
$[g(x)]_{i}\lambda_{i}=-\mu\,\Leftrightarrow\,s_{i}\lambda_{i}=\mu>0,$ (10b)
@@
Note that (10a) is derived from the original problem (1), not from the one with slack variables (9) (Nocedal and Wright, 2006). The parameter $\mu$ relaxes the complementary slackness and biases $s$ and $\lambda$ toward the feasible region, i.e., $s_{i},\lambda_{i}\geq 0$, which is known as the central path. The conditions (10) can also be obtained by adding constraints for $s$ to the objective, forming a modified problem
@@
When the condition above is satisfied, $\mu$ is reduced for the next iteration.
2.5 Primal Dual Augmented Lagrangian method
@@ 2.5.1 PDAL with slack variables
Consider the problem (9) again. The Primal Dual Augmented Lagrangian (PDAL) with slack variables is given by
@@
where $\lambda_{e}$ and $\nu_{e}$ are Lagrangian multiplier estimates for $\lambda$ and $\nu$ respectively. We take $\mu_{\rm{I}}=1/\rho_{\rm{I}}$ and $\mu_{\rm{E}}=1/\rho_{\rm{E}}$ as scalars for simplicity, but they can also be vectors of the corresponding size. Note that PDAL penalizes not only the violation of constraints but also the deviation of multipliers from the trajectory of the minimizers. Therefore, PDAL $\mathcal{L}_{\rm{PD}}$ is minimized for all variables, including Lagrangian multipliers, in contrast to simple AL, where the inner loop performs optimization with fixed multipliers. We now wish to minimize $L_{\rm{PD}}$ for $x,s,\lambda,\nu$. For this purpose, we first obtain the optimal $s$ denoted by $s^{\ast}$ for other variables. To keep the notation simple and compact, we only consider inequality constraints during the derivation of $s^{\ast}$, reducing $\mu_{\rm{I}}$ and $\mu_{\rm{E}}$ to $\mu$. Since the equality constraint terms are not affected by $s$, the complete form with inequalities is easily recovered after computing $s^{\ast}$. Completing the square of $\mathcal{L}_{\rm{PD}}$ in terms of $s$ yields
@@ 2.5.2 Minimization of PDAL
The matrix on the LHS becomes numerically unstable as $\mu$ in the denominators becomes small. This instability can be alleviated by the transformation given in the original work. Considering the Hessian of the Lagrangian $\nabla_{xx}\mathcal{L}=\nabla_{xx}f_{0}+\lambda_{e}g_{xx}+\nu_{e}h_{xx}$, and that of $\mathcal{L}_{\rm{PD}}$, $[2\pi_{\rm{I}}-\lambda]_{+}$ and $2\pi_{\rm{E}}-\nu$ can be seen as a new estimate of Lagrangian multipliers, which gives the update law of multipliers in the outer loop as
@@ 2.7 Sequential Quadratic Programming
$\nabla f_{0}(x)+H\delta x+[\nabla g(x)]^{\mathsf{T}}\lambda$ $=0,$ (26)

### v3

@@ Abstract
This paper provides an overview, analysis, and comparison of second-order dynamic optimization algorithms, i.e., constrained Differential Dynamic Programming (DDP) and Sequential Quadratic Programming (SQP). Although a variety of these algorithms have been proposed and used successfully, there exists a gap in understanding the key differences and advantages, which we aim to provide in this work. For constrained DDP, we choose methods that incorporate nonlinear programming techniques to handle state and control constraints, including Augmented Lagrangian (AL), Interior Point, Primal-Dual Augmented Lagrangian (PDAL), and Alternating Direction Method of Multipliers (ADMM). Both DDP and SQP are provided in single- and multiple-shooting formulations, where constraints that arise from dynamics are encoded implicitly and explicitly, respectively. As a byproduct of the review, we propose a single-shooting PDAL DDP that has more favorable properties than the standard AL variant, such as the robustness to the growth of penalty parameters. We perform extensive numerical experiments on a variety of systems with increasing complexity to investigate the quality of the solutions, the levels of constraint violation, and the sensitivity of final solutions with respect to initialization, as well as targets. The results show that single-shooting PDAL DDP and multiple-shooting SQP are the most robust methods. For multiple-shooting formulation, both DDP and SQP can enjoy informed initial guesses, while the latter appears to be more advantageous in complex systems. It is also worth highlighting that DDP provides favorable computational complexity and feedback gains as a byproduct of optimization as is.
@@ 1.1 Differential Dynamic Programming
Another approach to handling constraints is applying Nonlinear Programming (NLP) techniques to DDP. This includes the penalty barrier (Frisch, 1955; Fiacco and McCormick., 1968), Augmented Lagrangian (AL) (D., 1969; Hestenes, 1969), Interior Point (IP) (Byrd et al., 1999; Wächter and Biegler, 2005, 2006), Primal-Dual Augmented Lagrangian (PDAL) (Gill and Robinson, 2012; Robinson, 2007), and Alternating Direction Method of Multipliers (ADMM) (Boyd et al., 2011). We note that these techniques can be used to solve dynamic optimization for robotic applications, such as Vanroye et al. (2023).
The combination of DDP and the $\log$ barrier penalty method was proposed in Grandia et al. (2019), where the barrier function is relaxed to allow constraint violations. The penalty function’s coefficient, known as a penalty parameter, is driven to zero in the original formulation but is fixed in this DDP approach. Almubarak et al. (2022) uses a similar approach with an exact barrier function. Despite these approximations, they are shown to work sufficiently well in practice. AL DDP is the most widely used among these combinations of NLP with DDP for inequality constraints (Plancher et al., 2017; Pellegrini and Russell, 2020) and equality constraints (Kazdadi et al., 2021). The method is quite robust in terms of cost reduction, but it can violate constraints, especially in the early stage of optimization, where Lagrangian multipliers are inaccurate or penalty parameters are not large enough. Moreover, it can take many iterations to achieve strict feasibility. To alleviate this problem, the researchers proposed extensions in which algorithms switch from AL DDP to other methods when a trajectory approaches convergence (Lantoine and Russell, 2008; Howell et al., 2019; Aoyama et al., 2021). IP DDP (Pavlov et al., 2021) optimizes control variables, Lagrangian multipliers, and slack variables as the original IP method using DDP. PDAL DDP (Jallet et al., 2022a) was proposed most recently. This method is similar to the AL DDP, but the Lagrangian multipliers are also optimized using DDP in contrast to the AL variant.
Both AL and PDAL use penalty parameters that are increased during optimization. In AL, large penalty parameters are known to interrupt optimization, whereas PDAL is more robust to them as described in Robinson (2007). The PDAL DDP is presented in multiple-shooting formulation, which we elaborate on in the next paragraph. Using the Alternating Direction Method of Multipliers (ADMM), several variations of constrained DDP have been presented such as in Sindhwani et al. (2017); Zhou and Zhao (2020), which split the problem into smaller subproblems that are solved sequentially. Distributed ADMM-based DDP algorithms have also been proposed for handling constrained multi-agent control problems by utilizing the parallelizable nature of ADMM (Saravanos et al., 2023; Huang et al., 2023).
@@ 1.2 Sequential Quadratic Programming
SQP was introduced as a constrained optimization technique in Wilson (1963), showing its power in many fields including robotics (Yunt and Glocker, 2006, 2007; Yunt, 2011; Kuindersma et al., 2016; Posa et al., 2014), aerospace (Kenway and Martins, 2014; Kamyar and Taheri, 2014) and chemical engineering (Lucia et al., 2013). The approach relies on sequentially solving QP subproblems with quadratically approximated objectives and linearized constraints, generating new nominal trajectories. For dynamical systems, SQP encodes the constraints that arise from linearized dynamics at every time step as equality constraints. SQP also has single-shooting and multiple-shooting formulations. The multiple-shooting variant has state and controls as decision variables. The single-shooting variant has some variations. It can be achieved by solving the same subproblem as the multiple-shooting variant and propagating the state using the system’s dynamics. A technique called condensing (Diehl et al., 2002) can also be used to eliminate the state variables from the subproblems. There exists a significant amount of SQP variants that have been proposed in the literature recently for nonlinear optimal control and model predictive control (MPC) (Kouzoupis et al., 2018). One of the main difficulties of the SQP approach was its computational demands to solve QP subproblems. To alleviate this problem Pantoja and Mayne (1991) proposed a stage-wise version of SQP under control inequality constraints. In the same spirit of DDP, this approach solves smaller subproblems at every time step, rather than solving a problem for a whole time horizon. Another approach for improving the time complexity is to exploit the special sparse structure of the constraints arising from dynamics while solving QP. The growth of computational complexity with respect to time horizon can be reduced from cubic to linear, which is as good as DDP. This approach was proposed with active-set (Glad and Jonson, 1984), IP (Steinbach, 1994), and barrier (Wang and Boyd, 2010) methods. The same reduction is achieved by Riccati-based recursion for Linear Quadratic Regulator (LQR) problems (Dohrmann and Robinett, 1997; Rao et al., 1998; Jørgensen et al., 2004). Note that this technique is available only for multiple-shooting SQP because the sparse structure is lost by condensing (Diehl et al., 2009). It is worth noting that the recursion provides feedback gains as byproducts, as in DDP, although the gains have not been actively used. In the single-shooting variant, rather than working on the complexity, Singh et al. (2022) incorporates the DDP-style closed loop rollout, achieving a faster convergence speed. Recently, Jordana et al. (2023) proposed an approach similar to the multiple-shooting SQP, showing its superiority over single- and multiple-shooting DDPs (Mastalli et al., 2020) in tasks without state constraints. This work uses LQR and ADMM to solve QP subproblems efficiently and to enforce constraints. Inspired by Stellato et al. (2020), the matrices used in QP are not updated every iteration to improve computational efficiency.
1.3 Other Methods
The most straightforward approach for dynamic optimization is collocation methods, which discretize the problem in time and treat (possibly nonlinear) dynamics as equality constraints (Kelly, 2017). The discretization process and the time step, where equality constraints are applied, are known as transcription and collocation points, respectively. Any solver can be used to solve the transcribed problem. Since the method does not have any requirements for the representation of dynamics, such as an integration scheme, it is less restrictive than those we present in this paper. However, a naive implementation does not scale well with the size of the problem. This paper focuses on methods that can take advantage of the problem structure in dynamic optimization.
@@ 1.4 Contribution
Although DDP and SQP have a wealth of literature separately, there exist only few works comparing the two approaches. In the unconstrained case - where SQP coincides with Newton’s method with line search - the early works by Liao and Shoemaker (1992); Murray and Yakowitz (1984) had made a comparison of the two approaches. DDP was noted to have the same quadratic convergence properties as Newton’s method, but with the advantage of solving linear equations whose size remains constant w.r.t. the time horizon. However, given the variety of constrained DDP methods available, a thorough comparison of how these methods compare with SQP-based methods is still missing. A comparison between AL DDP and an SQP-based solver SNOPT (Gill et al., 2002) has been made in Howell et al. (2019), where SQP was shown to converge slower in wall clock time. Sindhwani et al. (2017), compared ADMM DDP with control-limited DDP Tassa et al. (2014) and SQP under only control constraints. Xie et al. (2017) also compared their constrained DDP and SQP (SNOPT) using state and control constrained robotic tasks under a time budget, showing the superiority of DDP. However, the details of SQP, including single- or multiple-shooting etc., are not presented.
To better understand the modern landscape of algorithms and how they relate to each other and the optimization literature, we compare these algorithms from derivations to performance in this paper. In addition, we propose single-shooting PDAL DDP that inherits the merit of PDAL over AL, that is, the robustness to the large penalty parameter, and add it to the comparison.
Our thorough numerical experiments with simple to complex dynamical systems reveal the difference of the algorithms in terms of the quality of solutions, the levels of constraint violation, iterations for convergence, and the sensitivity of the final solutions with respect to initialization. The results indicate that DDP frequently shows its capability to find better local minima, whereas SQP generally performs better in satisfying constraints. It is also shown that although both DDP and SQP have multiple-shooting formulations that can enjoy informed initial guesses, the SQP variant tends to work more reliably in complex systems. Furthermore, our analysis shows that single-shooting DDPs can offer lower computational complexity, especially when the system is underactuated and the problem has a long time horizon. This computational efficiency is one of the main motivations for users to choose DDP.
@@
Feedback Gain - d
@@
    Proposed in this paper.
@@
    When proper method is used as mentioned in section 1.2 and 5.2.
@@
    Although there exist multiple variants, here we use single-shooting with condensing. It cannot enjoy the reduction available in the multiple-shooting variant but can be reduced from $N^{3}m^{3}$ to $N^{2}m^{3}$ (Kouzoupis et al., 2018).
  • f
    Originally $N^{3}(n+m)^{3}$, but can be reduced to $N(n+m)^{3}$ with the sparse structure as mentioned in section 1.2 and 5.2.
@@ 2.2 Log Barrier methods
2.3 Augmented Lagrangian methods
@@
2.4 Interior Point methods
@@
$[g(x)]_{i}\lambda_{i}=0,\,\Leftrightarrow\,s_{i}\lambda_{i}=0$ (10b)
@@
Note that (10a) is derived from the original problem (1), not from the one with slack variables (9) (Nocedal and Wright, 2006). The complementary slackness (10b) is relaxed by the parameter $\mu$, i.e., $s_{i}\lambda_{i}=\mu(>0)$. The parameter biases $s$ and $\lambda$ toward the feasible region, i.e., $s_{i},\lambda_{i}\geq 0$, which is known as the central path. The relaxed condition of (10) can also be obtained by adding constraints for $s$ to the objective, forming a modified problem
@@
When the condition above is satisfied, $\mu$ is reduced for the next iteration, making the relaxed complementary slackness close to the actual one. There is an important mechanism implemented in IP-based packages such as IPOPT Wächter and Biegler (2006), which is known as feasibility restoration. When the line search cannot find an acceptable step size, even when the candidate step is smaller than a threshold, the feasibility restoration phase is invoked. In this phase, the algorithm focuses on minimizing infeasibility to find a solution that the filter can accept.
2.5 Primal-Dual Augmented Lagrangian method
@@ 2.5.1 PDAL with slack variables
Consider the problem (9) with constraints $h(x)=0$ as in (4). The Primal-Dual Augmented Lagrangian (PDAL) with slack variables is given by
@@
where $\lambda_{e}$ and $\nu_{e}$ are Lagrangian multiplier estimates for $\lambda$ and $\nu$ respectively. We take $\mu_{\rm{I}}=1/\rho_{\rm{I}}$ and $\mu_{\rm{E}}=1/\rho_{\rm{E}}$ as scalars for simplicity, but they can also be vectors of the corresponding size. Note that PDAL penalizes not only the violation of constraints but also the deviation of multipliers from the trajectory of the minimizers. Therefore, PDAL $\mathcal{L}_{\rm{PD}}$ is minimized for all variables, including Lagrangian multipliers, in contrast to simple AL, where the inner loop performs optimization with fixed multipliers. We now wish to minimize $\mathcal{L}_{\rm{PD}}$ for $x,s,\lambda,\nu$. For this purpose, we first obtain the optimal $s$ denoted by $s^{\ast}$ for other variables. To keep the notation simple and compact, we only consider inequality constraints during the derivation of $s^{\ast}$, reducing $\mu_{\rm{I}}$ and $\mu_{\rm{E}}$ to $\mu$. Since the equality constraint terms are not affected by $s$, the complete form with inequalities is easily recovered after computing $s^{\ast}$. Completing the square of $\mathcal{L}_{\rm{PD}}$ in terms of $s$ yields
@@ 2.5.2 Minimization of PDAL
The matrix on the LHS becomes numerically unstable as $\mu$ in the denominators becomes small. This instability can be alleviated by the transformation given in the original work Robinson (2007). Considering the Hessian of the Lagrangian $\nabla_{xx}\mathcal{L}=\nabla_{xx}f_{0}+\lambda_{e}g_{xx}+\nu_{e}h_{xx}$, and that of $\mathcal{L}_{\rm{PD}}$, $[2\pi_{\rm{I}}-\lambda]_{+}$ and $2\pi_{\rm{E}}-\nu$ can be seen as a new estimate of Lagrangian multipliers, which gives the update law of multipliers in the outer loop as
@@ 2.7 Sequential Quadratic Programming
$\nabla f_{0}(x)+[\nabla g(x)]^{\mathsf{T}}\lambda=0,$ (26)